Determine the concentration of a solution (in% by mass, % by volume, % by mass volume, molarity and ppm) containing 5 grams of NaCl and 50 ml of water. What is the mole fraction of NaCl in the solution?


1
Expert's answer
2021-04-07T11:37:51-0400

"\rho" (H2O)"=" 1 g/ml

m (H2O) "= \rho*V= 1 g/ml * 50ml = 50g"


1) % by mass = "\omega" (NaCl) "= m(NaCl)\/(m(H2O)+m(NaCl))*100%= 5g\/(5+50)g*100%" % = "5g\/(5+50)g*100%" % = 9.1%

2) % by mass volume = "m(NaCl)\/V""(H2O)" "*100%" % = 5g/50ml * 100% = 10 g/ml

3) n(NaCl) = m(NaCl)/M(NaCl) = 5g / 58.5g/mol = 0.0855 mol

molarity = M = n(NaCl)/V(H2O)= 0.0855 mol/ 0.05 L = 1.71 mol/L = 1.71 M

4) ppm = m(NaCl)/(m(NaCl)+m(H2O)) *106 = 5g/(50+5)g *106= 90909 ppm

5) n (H2O) = m(H2O) /M(H2O) = 50g / 18g/mol = 2.78 mol

mole fraction of NaCl ="\eta" (NaCl) = n(NaCl)/(n(H2O)+n(NaCl)) = 0.0855 mol / 2.78 mol = 0.03